Magnesium sulfate is often used in first

-

aid hot packs, giving off heat when

dissolved in water.

A

coffee

-

cup calorimeter at 25

o

C

contains 15.0 mL of water at

25

o

C

. A 2.00

-

g sample of MgSO

4

is dissolved in the water and 1.51 kJ of heat are

evolved. (You can make the following

assumptions about the solution: volume =

15.0 mL, density = 1.00 g/mL, specif

ic heat = 4.18

J/g

·

o

C

.)

a.

Write a balanced equation for the solution process.

b.

Is the process exothermic?

c.

What is

q

H

2

O

?

d.

What is the final temperature of the solution?

e.

What are the initial and final temperatures in

o

F?



1
Expert's answer
2021-11-16T12:24:02-0500

ΔHsolution = MCΔT


M = 15mL×1.00g/mL =15g


C = 4.18 J/g-degree Celsius


ΔT= Change in temperature = final temperature -initial temperature


ΔHsolution =1.51kJ=1.51×1000=1510J


Given, ΔHsolution = MCΔT


1510J = 15g×4.18 J/g-℃×ΔT


1510J = 62.7J/℃×ΔT


ΔT= 1510J÷62.7J/℃


= 24.08℃


ΔT= Change in temperature = final temperature -initial temperature


24.08℃ = final temperature-25 ℃


final temperature=24.08℃+25 ℃


=49.08℃


=49℃